The magnitude of vectorA - vectorB = A+B


Then find the angle between vectorA and vectorB?


waiting for reply

Grade:11

3 Answers

vikas askiitian expert
13 years ago

magnitude[A-B] = magnitude[A+B]

squaring both sides

A2 + B2 - 2A.B = A2 + B2 + 2A.B

 A.B = 0

A.B = ABCOS@ so

cos@ = 0

@ = pi/2

therefore angel bw A,B is 90o ...

 

approve if u like my ans

vijesh jayan
13 years ago

Thanks bro!!!!!!!!!!!!!!!!!!!!!

vikas askiitian expert
13 years ago

welcome my friend ..........